What is the NPV of the following project if the discount rate is 9%? Round to the nearest cent. Investment today: $-179,000; Cash flow in year 1: $54,000; Cash flow in year 2: $78,000; Cash flow in year 3: $54,000. ​[Hint: A negative cash flow is just an indication that it is an outflow. Don't enter it as a negative number in the NPV formula. Outflows should be factored in NPV calculations as positive numbers.
Answer #1

The cash flows are:
Year 0:$179000
Year 1:$54000
Year 2:$78000
Year 3:$54000

Net present value (NPV)= - Initial investment + Present value of future cash flows
Given that the discount rate is 9%.

Present value of future cash flows=54000/(1+9%)^1+78000/(1+9%)^2+54000/(1+9%)^3
=54000/1.09+78000/1.1881+54000/1.295029
=156890.2318

NPV=-179000+156890.2318=-$22109.7682